Let X be normally distributed with mean μ = 3,200 and standard deviation σ = 1,600. [Use Excel commands instead of the z table.] a. Find x such that P(X ≤ x) = 0.9382. (Round "z" value to 2 decimal places, and final answer to nearest whole number.) c. Find x such that P(3,200 ≤ X ≤ x) = 0.1217. (Round "z" value to 2 decimal places, and final answer to nearest whole number.)
Answer #1

@ (x + x) = 0.9982 - - P(x-7 P x 128000) = 0.9382 x - 3200 1600 ) = 0.9382 from standand z-table *-3200 = 1.54 =) a=5664 0 <for further any queries please comment and thank you